OpenStudy (anonymous):

is 1/3,2/3,1,4/3 arithmetic ?

OpenStudy (lgbasallote):

\[\frac 23 - \frac 13 = \frac 13 \\\\\\ \frac 14 - \frac23 = \frac{3}{12} - \frac{8}{12} = -\frac{5}{12}\] does that answer your questin?

OpenStudy (anonymous):

:s not really , it says if they are state the common difference

OpenStudy (lgbasallote):

oops lol sorry read the question wrong

OpenStudy (lgbasallote):

it was 1, 4/3 i thought it was 1/4

OpenStudy (lgbasallote):

\[\large \frac{2}{3} - \frac{1}{3} = \frac{1}{3}\] \[\large 1 - \frac{2}{3} = \frac{3}{3} - \frac 23 = \frac 13\] does that answer your question now?

OpenStudy (anonymous):

Basically, if you take each pair of numbers and subtract them, and the answer is always the same, there is a common difference. And if there is a common difference, then there is an arithmetic sequence.

OpenStudy (anonymous):

thanks guys ! im taking these quizes on my online class so i can graduate this friday. :D

OpenStudy (lgbasallote):

i wish my graduation depended on solving fractions :(

OpenStudy (anonymous):

my whole online class if just algebra :( not my strongest